# Near-linear Time Dispersion of Mobile Agents

### v2

@@ 14                   $a_{i}$ moves to $u_{i}$ if $a_{i}$ finds $\psi(u_{i})$ at $u_{i}$ then
The unique leader $a_{\mathrm{max}}$ invokes $\mathtt{Probe}()$ only when it settles an agent, except for when $a_{\mathrm{max}}$ itself becomes settled. Therefore, $a_{\mathrm{max}}$ invokes $\mathtt{Probe}$ exactly $k-1$ times, with precisely $k-i$ explorers present at the $i$-th invocation. By Lemma 7, the total number of steps required for the $k-1$ executions of $\mathtt{Probe}()$ is at most
$\sum_{\ell=1}^{k-1}(\log k-\log\ell)+O(k)\leq k\log k-\log(k!)+\log k+O(k)=O(k),$
@@
where we apply Stirling’s formula, i.e.$\log(k!)=k\log k-k+O(\log k)$. As demonstrated in Section 3, both forward and backward moves also require a total time of $O(k)$. Thus, $\mathbf{RootedOptimal}$ completes in $O(k)$ time. Regarding space complexity, the array variable $\mathtt{checked}$ is the primary factor, needing $O(\delta)$ bits per agent. Other variables require only $O(\log\delta)$ bits. ∎
6 Conclusion
@@
In this paper, we introduced a novel technique HEO-DFS and presented near time-optimal algorithms both for the rooted setting and the general setting.
A drawback of our HEO-DFS is that it essentially requires a synchronous model even in the rooted setting, i.e., it does not work on an asynchronous model. In HEO-DFS, we try to find an unvisited neighbor of the current location with the help of agents settled on other neighbors. These agents must return to their homes once the probing process finishes. In an asynchronous model, the unsettled agents may visit those homes before their owners return, disrupting the consistent behavior of HEO-DFS.
It remains open whether there is a $o(m_{k})$-time algorithm that solves the dispersion problem in the asynchronous model.

### v3

@@ 5 For Further Improvement in Time Complexities
The unique leader $a_{\mathrm{max}}$ invokes $\mathtt{Probe}()$ only when it settles an agent, except for when $a_{\mathrm{max}}$ itself becomes settled. Therefore, $a_{\mathrm{max}}$ invokes $\mathtt{Probe}$ exactly $k-1$ times, with precisely $k-i$ explorers present at the $i$-th invocation. By Lemma 7, the total number of steps required for the $k-1$ executions of $\mathtt{Probe}()$ is at most $\sum_{\ell=1}^{k-1}(\log k-\log\ell+O(1))=k\log k-(\log(k!)-\log k)+O(k)=O(k)$, where we apply Stirling’s formula, i.e.$\log(k!)=k\log k-k+O(\log k)$. As demonstrated in Section 3, both forward and backward moves also require a total time of $O(k)$. Thus, $\mathbf{RootedOpt}$ completes in $O(k)$ time. Regarding space complexity, the array variable $\mathtt{checked}$ is the primary factor, needing $O(\Delta)$ bits per agent. Other variables require only $O(\log\Delta)$ bits. ∎
6 Concluding Remarks
@@
In this paper, we introduced a novel technique HEO-DFS and presented near time-optimal dispersion algorithms based on this techniques both for the rooted setting and the general setting. Both the algorithms require only $O(\log(k+\Delta)$ bits per agent. Moreover, we give time-optimal algorithms that solves dispersion in the rooted setting at the cost of increasing the memory space to $O(\Delta+\log k)$ bits per agent.
It is worth mentioning that while HEO-DFS does not function in a fully asynchronous model, where movement between two nodes may require an unbounded period, it does not require a fully synchronous model in the rooted setting. Specifically, $\mathbf{RootedDisp}$ and $\mathbf{RootedOpt}$ can operate under an asynchronous scheduler if every movement of agents between nodes is atomic, i.e., each agent is always located at a node and never on an edge at any time step. Under this scheduler, after the probing process is completed at a node $v$, unsettled agents can wait for all helping settlers to leave $v$ before they themselves depart. Since every movement is atomic, when unsettled agents visit the home node of one of these settlers, the settler has already returned. Thus, $\mathbf{RootedDisp}$ and $\mathbf{RootedOpt}$ functions under any fair scheduler that guarantees every movement is atomic. However, this move-atomicity is not sufficient for $\mathbf{GeneralDisp}$, because this algorithm, designed for the general setting, differentiates the moving speeds of agents based on their roles – leader, strong zombie, or weak zombie – which inherently requires a fully synchronous scheduler.
Appendix A How to Terminate $\mathbf{RootedDisp}$ and $\mathbf{RootedOpt}$
The termination can be easily implemented in a common way both for $\mathbf{RootedDisp}$ and $\mathbf{RootedOpt}$.
The last settled agent $a_{\mathrm{last}}$ recognizes the achievement of dispersion immediately after all agents are settled at distinct nodes. In $\mathbf{RootedOpt}$, the last settled agent $a_{\mathrm{last}}$ may not be uniquely determined when the unsettled agents at node $w$ detect that $|U_{0}(w)|\geq\ell$, where $\ell$ is the number of those agents, and they move to and settle at $\ell$ distinct nodes (line 33 in Algorithm 6). In this case, those $\ell$ agents immediately terminate and are not used for the subsequent termination detection. Instead, we define $a_{\mathrm{last}}$ as $\psi(w)$, who settled at $w$ before those $\ell$ agents.
Roughly speaking, $a_{\mathrm{last}}$ then relays the fact that dispersion is achieved to all remaining agents via edges in the DFS tree, say $T$, that the agents have constructed so far. For any settled node $v\neq w=\nu(a_{\mathrm{max}})$, we define $f(v)$ as the next node of $v$ in the path leading to $w$ in tree $T$. Formally, we define $f(v)=u_{1}$, where $p=u_{0},u_{1},\dots,u_{s}\ (v=u_{0},\ w=u_{s})$ is the unique path between $v$ and $w$. Like the time slots in $\mathbf{GeneralDisp}$, we divide time steps into three time slots: 1, 2, and 3. Specifically, we classify time step $t$ as slot $i\in\{1,2,3\}$ if $(t\bmod 3)+1=i$. In every time slot 1, each agent moves or stays according to the original algorithms, i.e., Algorithms 1 and 6 for $\mathbf{RootedDisp}$ and $\mathbf{RootedOpt}$, respectively. In slot 2, each settled agent $b$ with home $\nu(b)$ moves to $f(\nu(b))$, and in slot 3, it goes back to $\nu(b)$. In slots 2 and 3, no unsettled agent moves. After $a_{\mathrm{last}}$ detects the achievement of dispersion at time step $t$, which must be slot 1, it does not terminate immediately. It waits for other settlers coming to $w$ in the next slot 2, tells them about the achievement of dispersion in slot 3, and halts in the next slot 1. The settlers who newly learn this fact go back to their homes in slot 3 and execute the same procedure: wait until the next slot 3, inform the arriving settlers about the achievement, and halt in the next slot 1. Since the diameter of the DFS tree $T$ is at most $k-1$, this termination procedure only requires $O(k)$ time.
Each settler $b$ with home $v=\nu(b)$ must recognize which ports $10,1,\dots,\delta_{v}-$ 1 lead to node $f(v)$. This recognition can be easily implemented. For $\mathbf{RootedDisp}$, settler $b$ knows $f(v)=N(v,b.\mathtt{next})$ if $b.\mathtt{next}\neq\bot$, and $f(v)=N(v,b.\mathtt{parent})$ otherwise. We slightly modify $\mathbf{RootedOpt}$ for this recognition: (i) settler $b$ maintains a variable $b.\mathtt{next}$, (ii) $b$ sets $b.\mathtt{next}\leftarrow u$ in Line 6, and (iii) $b$ sets $b.\mathtt{next}\leftarrow b.\mathtt{parent}$ in Line 11. Then, $b$ always knows $f(v)=N(v,b.\mathtt{next})$.